unicornicecream93:

xpression is shown below f(x)=2x^2+5x-3 Part A: What are the x-intercepts of the graph of f(x)? Show your work factoring and using the Zero Product Property. Part B: Is the vertex of the graph of f(x) going to be a maximum or minimum? What are the coordinates of the vertex? Justify your answers and show your work using the vertex formula. Part C: Identify the point of the y-intercept for f(x). Part D: What are the steps you would use to graph f(x)? Justify that you can use the answers obtained in Part A, Part B, and Part C to draw the graph. Draw the graph

snowflake0531:

it wouldbe easier with a screenshot

unicornicecream93:

Here it is

100kpimpcess:

@Laylalyssa

unicornicecream93:

Urgent due by end of day

100kpimpcess:

@supie

jhonyy9:

f(x)= 2x^2 +5x -3 = (x -1/2)(x +3) =(2x -1)(x+3) questions ?

hotboy3245657868760:

f(x)= 2x^2 +5x -3 = (x -1/2)(x +3) =(2x -1)(x+3)

unicornicecream93:

Ugh help

snowflake0531:

just.. they already wrote (2x-1)(x+3) take it apart 2x-1=0, so x=1/2 x+3=0, so x=-3 so the x-intercepts are at 1/2 and -3

unicornicecream93:

Part b part and part d

BigE21:

have you ever used desmos? That could help it helped me *desmos graphing calculator*

unicornicecream93:

How, I just need answers it's due by the end of the day

BigE21:

search Desmos graphing calculator and you type in the equation dont forget to put in f(x) in the beginning and it will graph the equation on a graph for you might not be able to help specifically on this question but it does on a lot of them

unicornicecream93:

What's the y intercepts though

unicornicecream93:

Umm what's the vertex coordinates

jhonyy9:

@supie opinion pls. ? ty.